OpenStudy (anonymous):

Find two consecutive odd numbers such that the sum of three-sevenths of the first number and one-third of the second number is equal to thirty-eight.

OpenStudy (anonymous):

49 and 51

OpenStudy (anonymous):

Let x = the smaller number, then\[\text{Solve}\left[\left(\frac{3}{7}\right)x+\left(\frac{1}{3}\right)(x+2)=38,x\right] \]x = 49